OpenStudy (anonymous):

1. f(x) = 3 - x2 - 6x 2. (x) = x2 - 8x + 2 I have these two problems. I need to find out the domain, range, maximum, and minimum for both of the above. I keep getting the wrong answers when I try it, and some that don't make sense.. Please help? Thanks in advance!

OpenStudy (anonymous):

Can someone help me please!!?

OpenStudy (solomonzelman):

For a parabola with a positive leading coefficient, absolute maximum won't exit. (it opens up, and its absolute minimum is the vertex - y-coordinate is the value of the minimum, and x-coordintae is where this minimum is located) ---------------------------------------------------- For a parabola with a negative leading coefficient, absolute minimum won't exit. (it opens up, and its absolute maximum is the vertex - y-coordinate is the value of the maximum, and x-coordintae is where this maximum is located) ---------------------------------------------------- To find the vertex in each parabola, you need to complete the square. (for each function)

OpenStudy (solomonzelman):

``` If you are not familiar with what "perfect square trinomial" means, then I would advise to review that concept (here, with other people, or watch a video, read a book, get a tutor.... idk, that is your responsibility. I won't do it now, because I got to go pretty soon). ```

OpenStudy (jtvatsim):

@macky342 You still there? Have you made any progress on the question, or still stuck? :)

OpenStudy (anonymous):

I'm still stuck unfortunately.

OpenStudy (jtvatsim):

K, let's see if we can get somewhere...

OpenStudy (anonymous):

math has always been my worst subject. thank you.

OpenStudy (jtvatsim):

Well, maybe we'll be able to turn that around. :) I'm taking a look...

OpenStudy (anonymous):

Great! Thank you. This is the last question i have in this class before i graduate so i just want to be done!

OpenStudy (jtvatsim):

First thing I'm going to do is rewrite the equations this way: \[f(x) = -x^2-6x+3\] and \[g(x) = x^2-8x+2\].

OpenStudy (anonymous):

okay, now what do i do?

OpenStudy (jtvatsim):

Alright, personally, for me a picture is worth a thousand words, so I'm going to set these up to make it easier to graph a picture.

OpenStudy (jtvatsim):

This is a trick that most classes don't teach you, but I'm going to factor an x out of the first two terms of each. The result looks like this:

OpenStudy (jtvatsim):

\[f(x)=-x(x+6)+3\] and \[g(x)=x(x-8)+2\]

OpenStudy (jtvatsim):

You probably haven't seen that, but does that make sense so far?

OpenStudy (anonymous):

not really :(

OpenStudy (jtvatsim):

OK, no worries, all I did is this for the first one \[ f(x) = -x^2-6x+3 = x(-x-6)+3 = -x(x+6)+3 \] So, I just ignored the 3, and factored an x out of the first two terms. Then, I took out the negative sign (I didn't have to but it looks nicer). Please let me know if you have any questions on this part. We may need to review factoring. :)

OpenStudy (anonymous):

i am not understanding a thing thats going on. i reviewed the chapter over 5 times. had a tutor try and explain it and my last hope was this website and i feel like theres still no hope :(

OpenStudy (jtvatsim):

Gotcha. No worries, there's just a mental block somewhere that we have to sort out. I've had plenty before myself. :) Let's start more basic.

OpenStudy (anonymous):

perfect. thanks.

OpenStudy (jtvatsim):

I'm going to do a few diagnostic questions (totally different than the current question) to see where your understanding starts. Does this make sense? \[x^2 + x = x(x + 1)\]

OpenStudy (anonymous):

nope.

OpenStudy (jtvatsim):

OK, that's fine. Let's go simpler. :)

OpenStudy (anonymous):

the x's is where i get lost.

OpenStudy (anonymous):

I'm pretty good at basic math

OpenStudy (jtvatsim):

OK, so does this make more sense \[3^2 + 3 = 3(3+1)\]

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, so let's see if we can build on that understanding.... Let me think for a moment...

OpenStudy (mertsj):

Do you know that these two equations have graphs that are parabolas?

OpenStudy (anonymous):

yes.

OpenStudy (mertsj):

In the first one, the x^2 term has a negative coefficient so it will look like this:

OpenStudy (mertsj):



OpenStudy (anonymous):

okay.

OpenStudy (mertsj):

In the second one, the x^2 term is positive so it will look like this:

OpenStudy (mertsj):